Kolby’s Korndogs is looking at a new sausage system with an installed cost of $715,000. This cost will be depreciated straight-line to zero over the project’s 6-year life, at the end of which the sausage system can be scrapped for $97,000. The sausage system will save the firm $207,000 per year in pretax operating costs, and the system requires an initial investment in net working capital of $59,000. If the tax rate is 22 percent and the discount rate is 8 percent, what is the NPV of this project?
Answer #1

Answer 1  
Year 0 Cash flow  
  
Cost of machine   -$715,000
Increase in net

working capital   -59000
___________________________  
Total Cash flow at year 0   -$774,000
___________________________  

Cash flow from Year 1 to 6

Year 1-6

Savings in cost 207000.00

less: Depreciation -119166.67

(715000/6)

___________________________

Operating profit 87833.33

less tax @ 22% -19323.33

___________________________

Profit after tax 68510.00

Add: Depreciation 119166.67

___________________________

Free cash flow 187676.67

Note: Depreciation is non cash expenses. It is only deducted for tax calculation. Again it will be added as it is not in cash.

Year 6 Terminal cash inflows   

Salvage value.    97000
Less: Tax on sale.   -21340
Add: Working capital

recovered.    59,000.00
___________________________  
Terminal cash inflow   134,660.00
___________________________  
Note: Book value of asset after 6 year Depreciation is 0   0.00
Sale value is   97000
So Capital gain = sale - Book value

=97000-0=   97,000.00
Tax @22% paid on gain=   21340


NpV = sum of all free cash flows          
PVF formula = 1/(1+i)^n          
for year 1 PVF =1/(1+8%)^1=   0.9259259259      
for year 2 PVF =1/(1+8%)^2=   0.8573388203   and So on.  
          
Calcultion of NPV          
Year   Free cash flow   PVF @8%   PV = Cash flow * PVF
0   -$774,000   1.0000   -774,000.00
1   187676.67   0.9259   173,774.69
2   187676.6667   0.8573   160,902.49
3   187676.6667   0.7938   148,983.79
4   187676.6667   0.7350   137,947.95
5   187676.6667   0.6806   127,729.59
6 322,336.67   0.6302   203,126.78
          
NpV = sum of all free cash flows=           178,465.29
          
So NPV of project is    178,465.29         

Note : year 6 cash inflow includes FCF and Terminal cash inflows.